A pair of fair dice is tossed one time. What is the probability that ⦁ The sum of two faces is a prime number? ⦁ The sum of two faces is more than 9? ⦁ The sum of two faces is between 5 and 9? ⦁ The sum of two faces is at most 6? ⦁ The sum of two faces is at least 8?


1
Expert's answer
2021-06-01T03:44:33-0400


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
1 & 2 & 3 & 4 & 5 & 6 & 7 \\
 \hdashline
2 & 3 & 4 & 5 & 6 & 7 & 8\\
 \hdashline
3 & 4 & 5 & 6 & 7 & 8 & 9 \\
 \hdashline
4 & 5 & 6 & 7 & 8 & 9 & 10 \\
 \hdashline
5 & 6 & 7 & 8 & 9 & 10 & 11 \\
 \hdashline
6 & 7 & 8 & 9 & 10 & 11 & 12 \\
 \hdashline
\end{array}"

1.

Prime numbers: "2,3,5,7,11"

"P(sum\ is\ prime)=\dfrac{1+2+4+6+2}{36}=\dfrac{5}{12}"



2.


"P(sum>9)=\dfrac{3+2+1}{36}=\dfrac{1}{6}"

3.


"P(5\leq sum\leq 9)=\dfrac{4+5+6+5+4}{36}=\dfrac{2}{3}"

"P(5< sum<9)=\dfrac{5+6+5}{36}=\dfrac{4}{9}"

4.


"P( sum\leq 6)=\dfrac{1+2+3+4+5}{36}=\dfrac{5}{12}"

5.

"P(sum\geq 8)=\dfrac{5+4+3+2+1}{36}=\dfrac{5}{12}"